Randomly selected birth records were​ obtained, and categorized as listed in the table to the right. Use a

0.01

0.01 significance level to test the reasonable claim that births occur with equal frequency on the different days of the week. How might the apparent lower frequencies on Saturday and Sunday be​ explained?

Day nbsp

Day

Sun

Mon

Tues

Wed

Thurs

Fri

Sat

Number of Births

Number of Births

48

48

64

64

61

61

60

60

56

56

65

65

47

47

Determine the null and alternative hypotheses.

H0​: (choose one)

  1. At least one day has a different frequency of births than the other days.
  2. Births occur with the same frequency on the different days of the week.
  3. Births occur with all different frequencies on the different days of the week.
  4. At least two days have a different frequency of births than the other days.

H1​: (choose one)

  1. At least one day has a different frequency of births than the other days.
  2. Births occur with the same frequency on the different days of the week.
  3. Births occur with all different frequencies on the different days of the week.
  4. At least two days have a different frequency of births than the other days.

Calculate the test​ statistic, χ2 = _____

​(Round to three decimal places as​ needed.)

Calculate the​ P-value.

​P-value = _____

​(Round to four decimal places as​ needed.)

What is the conclusion for this hypothesis​ test?

A. Fail to reject H0. There is sufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

B.

Fail to reject H0. There is insufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

C.

Reject H0. There is sufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

D.

Reject H0. There is insufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

How might the apparent lower frequencies on Saturday and Sunday be​ explained?

A.

Induced or​ Caesarean-section births are scheduled during the week whenever possible.

B.

Induced or​ Caesarean-section births are scheduled on weekends whenever possible.
Answer #1

Null Hypotheses, H0​:

Births occur with the same frequency on the different days of the week.

Alternative Hypothesis, H1:

At least one day has a different frequency of births than the other days.

Day O: Observed E: Expected (O-E)2/E
Sun 48 401/7 =57.28571429 1.505165657
Mon 64 57.28571429 0.786961168
Tues 61 57.28571429 0.240826505
Wed 60 57.28571429 0.128607054
Thurs 56 57.28571429 0.02885643
Fri 65 57.28571429 1.038831493
Sat 47 57.28571429 1.846811542
Total 401 401 5.576

\chi^2 =\Sigma \ [(O-E)^2/E] =5.576

P-value for Chi-square test statistic of 5.576 at n-1 =7-1 =6 degrees of freedom is: P-value =0.4723

Conclusion: Since, P-value is greater than the significance level of 0.01 (i.e., 0.4723 > 0.01), we fail to reject the null hypothesis(H0) at 0.01 significance level.

Thus, the correct option is B. Fail to reject H0. There is insufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

The apparent lower frequencies on Saturday and Sunday might be explained as:

B.

Induced or​ Caesarean-section births are scheduled on weekends whenever possible.

Therefore, the number of births on Saturdays and Sundays are less as the Caesarean-section births are less common (about 1 in 3) and the number of births during the other days of the week are relatively more.